The color lines of hydrogen are due to the transitions from higher-energy state, n high, to the lower- energy state (n 2). UsNot sure how to solve this..
Answer #1

Solut theory for the from the Rydbergs hydrogen spectra 17 wavelength is given by equation } = Rp (Thin RH - Rydbergs conet487.32 becnaugyz = traz °C Minya) = 7,20 ... (Perign) 2 = 4x7.32 (Monigh) = 18.81 S hnign = 2.96 ... 3.32 nhigh = 3 - 2020.0812=2 put values 447.18109 x 1.0977107 - (2 (2) ? x 442.1715 X 1.097 7107 Chmiguja > 4.90-4 - 4X4.90 (nnigh) - 4x4.90 0.90 (hn(nnign) = 1 21. 3 29.28 einnigh = 3.12 [nnigh a 3 color Red Brue violet I wavelength (nm) - 667.6 4471 667.6 | Assignment is